40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 21:16:32 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
    Reply  Tue Mar 3 19:26:44 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
       Reply  Tue Mar 3 23:59:26 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
          Reply  Wed Mar 4 03:37:51 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
Message ID: 1352     Entry time: Wed Mar 4 03:37:51 2009     In reply to: 1351
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Low-gain High-gain PD switching may not be working well 

Quote:

This sounds like the QPD whitening gain sliders may be stuck. The slider twiddling script should be run, or the sliders should be twiddled by hand.


Yes, it was indeed the whitening gain slider problem.
I moved them and the QPDX gain went up suddenly. I reinstalled the BS in front of the QPDX and adjusted the offsets, gains accordingly.
Now the high-gain to low-gain switching is fine.
ELOG V3.1.3-